OpenStudy (anonymous):

Please help! Write the standard form equation of a hyperbola with foci at (0,8) and (0,-8) and eccentricity at 4/3

OpenStudy (anonymous):

so far

OpenStudy (anonymous):

i did 0+0/2 and -8+8/2 to get the center at (0,0)

OpenStudy (anonymous):

\[y ^{2}-x ^{2}=1\]

OpenStudy (anonymous):

if e=c/a then c^2=16 and a^2=9

OpenStudy (anonymous):

if i use b^2=c^2-a^2

OpenStudy (anonymous):

\[b=\sqrt{7} \] and b^2=7

OpenStudy (anonymous):

@ganeshie8 @ShailKumar

OpenStudy (anonymous):

do you think its correct @ShailKumar?

OpenStudy (anonymous):

if e=c/a then c^2=16 and a^2=9. How ?

OpenStudy (anonymous):

because e=4/3 if e=c/a then c=4 and c=3 c^2=16 and a^2=9

OpenStudy (anonymous):

sorry * a=3

OpenStudy (anonymous):

But I can say that c = 8 and a = 6 !

OpenStudy (anonymous):

4^2=16 and 3^2=9 though because in standard form we want (x-h)^2 over a^2 and (y-k)^2 over b^2

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

In e = c/a, you know c = 4/3. but you have two unknowns a and c.

OpenStudy (anonymous):

For two unknowns, you need two equations ! right ?

OpenStudy (anonymous):

is the center (0,0) correct? right sorry we just solved one like this!

OpenStudy (anonymous):

Center is correct.

OpenStudy (anonymous):

We need one more equation two solve for a and c.

OpenStudy (anonymous):

What is distance of the center from any of the foci ?

OpenStudy (anonymous):

distance between x or y?

OpenStudy (anonymous):

oh -8 and 8 duh

OpenStudy (anonymous):

-8-0=-8 8-0=8

OpenStudy (anonymous):

Right ! Now , the hyperbola is a vertical one.

OpenStudy (anonymous):

yep! so its going to be (y-k)^2/a^2-(x-h)^2/b^2 because its parallel to y-axis

OpenStudy (anonymous):

Yup, so, c = 8 and since e = c/a ==> a = 6

OpenStudy (anonymous):

Find b

OpenStudy (anonymous):

how is a=6 again?

OpenStudy (anonymous):

b=square root 28 and square root 28 squared is just 28 so b^2=28

OpenStudy (anonymous):

Right! Do you understand why a = 6 ?

OpenStudy (anonymous):

You got it right.

OpenStudy (anonymous):

yay! but no, why is a=6?

OpenStudy (anonymous):

e = 4/3, c = 8 , e = c/a so a = 6

OpenStudy (anonymous):

so you just multiplied 4x2=8 c=8 and 3x2=6 so a=6?

OpenStudy (anonymous):

the way i see it...

OpenStudy (anonymous):

Ok, do you understand why c = 8 ?

OpenStudy (anonymous):

same thing, you multiplied 4x2=8

OpenStudy (anonymous):

wait, no c=8 because its the distance from foci to center

OpenStudy (anonymous):

got it, thanks!!!

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Ok, signing off.. :)